OpenStudy (anonymous):

Divide. 118.23 ÷ 2.1

OpenStudy (kirbykirby):

The best trick for decimal division is to keep multiplying your numbers by ten till you don't have any decimal point: 118.23*10 ÷ 2.1* 10 = 1182.3 ÷ 21 again: 1182.3*10 ÷ 21*10 = 11823 ÷ 210. Now you just perform long division

OpenStudy (anonymous):

thanks!

OpenStudy (kirbykirby):

:)

OpenStudy (anonymous):

:)